# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

### v2

@@ 4.2 Fusion Mechanisms
4.2.1 Score Fusion and Alignment
Models in this category utilize distinct strategies to align multimodal representations. Zhi Lim et al. (2024) convert text, tables, and images into a single textual format using a cross-encoder trained for relevance scoring. Sharifymoghaddam et al. (2024) introduce interleaved image–text pairs that vertically merge multiple few-shot images (as in LLaVA Liu et al. (2023a)), while aligning modalities via CLIP score fusion Hessel et al. (2021) and BLIP feature fusion Li et al. (2022a). Riedler and Langer (2024), Wiki-LLaVA Caffagni et al. (2024), C3Net Zhang et al. (2024c), and MegaPairs Zhou et al. (2024a), embed images and queries into a shared CLIP space.
@@ 4.2.1 Score Fusion and Alignment
VISA Ma et al. (2024b) employs the Document Screenshot Embedding (DSE) model to align textual queries with visual document representations by encoding both into a shared embedding space. REVEAL Hu et al. (2023) injects retrieval scores into attention layers to minimize L2-norm differences between query and knowledge embeddings, and MA-LMM He et al. (2024) aligns video-text embeddings via a BLIP-inspired Query Transformer Li et al. (2022a). LLM-RA Jian et al. (2024) concatenates text and visual embeddings into joint queries to reduce retrieval noise, while RA-BLIP Ding et al. (2024b) employs a 3-layer BERT-based adaptive fusion module to unify visual–textual semantics. Xue et al. (2024) use a prototype-based embedding network Zheng et al. (2023) to map object-predicate pairs into a shared semantic space, aligning visual features with textual prototypes. Re-IMAGEN Chen et al. (2022b) balances creativity and entity fidelity in text-to-image synthesis via interleaved classifier-free guidance during diffusion sampling. To improve multimodal alignment, VISRAG Yu et al. (2024) enhances alignment with position-weighted mean pooling on VLM hidden states, prioritizing later tokens for relevance, and RAG-Driver Yuan et al. (2024) aligns visual–language embeddings using visual instruction tuning and an MLP projector.
4.2.2 Attention-Based Mechanisms
Attention-based methods dynamically weight cross-modal interactions to support task-specific reasoning. EMERGE Zhu et al. (2024b), MORE Cui et al. (2024), and AlzheimerRAG Lahiri and Hu (2024) integrate heterogeneous data via cross-attention. RAMM Yuan et al. (2023) employs a dual-stream co-attention transformer, combining self-attention and cross-attention to fuse retrieved biomedical images/texts with input data. RAGTrans Cheng et al. (2024) applies user-aware attention to social media features.
@@ 4.2.2 Attention-Based Mechanisms
For video-text alignment, MV-Adapter Jin et al. (2024) leverages Cross Modality Tying to align embeddings, and M2-RAAP Dong et al. (2024b) enhances fusion through an auxiliary caption-guided strategy that re-weights frames and text captions based on intra-modal similarity. A mutual-guided alignment head then filters misaligned features using dot-product similarity and frame-to-token attention, generating refined frame-specific text representations. Xu et al. (2024a) condition text generation on visual features using gated cross-attention, and Mu-RAG Chen et al. (2022a) employs intermediate cross-attention for open-domain QA. Kim et al. (2024) leverage cross-modal memory retrieval with pre-trained CLIP ViT-L/14 to map video-text pairs into a shared space, enabling dense captioning through the attention-based fusion of retrieved memories.
4.2.3 Unified Frameworks and Projections
Unified frameworks and projection methods consolidate multimodal inputs into coherent representations. Su et al. (2024a) employ hierarchical cross-chains and late fusion for healthcare data, while IRAMIG Liu et al. (2024b) iteratively integrates multimodal results into unified knowledge representations. M3DocRAG Cho et al. (2024) flattens multi-page documents into a single embedding tensor, and PDF-MVQA Ding et al. (2024c) fuses Region-of-Interest (RoI)-based and patch-based (CLIP) vision-language models Long et al. (2022).
@@ 4.2.3 Unified Frameworks and Projections
DQU-CIR Wen et al. (2024) unifies raw data by converting images into text captions for complex queries and overlaying text onto images for simple ones, then fusing embeddings via MLP-learned weights. SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input into unimodal text for subsequent processing. UFineBench Zuo et al. (2024) utilizes a shared granularity decoder for ultra-fine text–person retrieval. Nguyen et al. (2024) introduce Dense2Sparse projection, converting dense embeddings from models like BLIP/ALBEF Li et al. (2022a) into sparse lexical vectors using layer normalization and probabilistic expansion control to optimize storage and interpretability.
@@ 4.3 Augmentation Techniques
Basic RAG systems typically follow a single retrieval step and pass retrieved content directly to the generation phase without further refinement, which can lead to inefficiencies and suboptimal outputs. To address this, augmentation techniques have been introduced to refine retrieved data beforehand, improving multimodal interpretation, structuring, and integration Gao et al. (2023).
4.3.1 Context Enrichment
@@
This focuses on enhancing the relevance of retrieved knowledge by refining or expanding retrieved data. General approaches incorporate additional contextual elements (e.g., text chunks, image tokens, structured data) to provide a richer grounding for generation Caffagni et al. (2024); Xue et al. (2024). EMERGE Zhu et al. (2024b) enriches context by integrating entity relationships and semantic descriptions. MiRAG Adjali et al. (2024) expands initial queries through entity retrieval and reformulation, enhancing subsequent stages for the visual question-answering. Video-RAG Luo et al. (2024b) enhances long-video understanding through Query Decoupling, which reformulates user queries into structured retrieval requests to extract auxiliary multimodal context. Img2Loc Zhou et al. (2024e) enhances accuracy by including both the most similar and the most dissimilar points from the database in the prompt, allowing the model to rule out implausible locations for its predictions.
4.3.2 Adaptive and Iterative Retrieval
@@
For more complex queries, dynamic retrieval mechanisms have proven effective. Adaptive retrieval approaches optimize relevance by adjusting retrieval dynamically. SKURG Yang et al. (2023) determines the number of retrieval hops based on query complexity. SAM-RAG Zhai (2024) and mR2AG Zhang et al. (2024f) dynamically assess the need for external knowledge and filter irrelevant content using MLLMs to retain only task-critical information. MMed-RAG Xia et al. (2024a) further improves retrieval precision by discarding low-relevance results, while OmniSearch Li et al. (2024b) introduces a self-adaptive retrieval agent that dynamically decomposes complex multimodal questions into sub-question chains and plans retrieval actions in real-time based on retrieved content.
Iterative approaches refine results over multiple steps by incorporating feedback from prior iterations. IRAMIG Liu et al. (2024b) improves multimodal retrieval by dynamically updating queries based on retrieved content. Similarly, OMG-QA Nan et al. (2024) introduces a multi-round retrieval strategy, where each retrieval step incorporates episodic memory to refine subsequent queries. An evaluation module assesses retrieval effectiveness at each step, guiding the refinement of subsequent retrieval efforts through feedback. RAGAR Khaliq et al. (2024) further enhances contextual consistency by iteratively adjusting retrieval based on prior responses and multimodal analysis.
@@ 4.4 Generation Techniques
Recent advancements in multimodal RAG generation focus on robustness, cross-modal coherence, and task-specific adaptability. These innovations can be broadly categorized into these sections:
4.4.1 In-Context Learning
@@
In-context learning (ICL) with retrieval augmentation enhances reasoning in multimodal RAGs by leveraging retrieved content as few-shot examples without requiring retraining. Models such as RMR Tan et al. (2024), Sharifymoghaddam et al. (2024), and RA-CM3 Yasunaga et al. (2023), extend this paradigm to multimodal RAG settings.
RAG-Driver Yuan et al. (2024) refines ICL by retrieving relevant driving experiences from a memory database, ensuring scenario-specific contextual alignment. MSIER Luo et al. (2024a) further enhances example selection through a Multimodal Supervised In-Context Examples Retrieval framework, leveraging a foundation MLLM scorer to evaluate both textual and visual relevance. Meanwhile, Raven Rao et al. (2024) introduces Fusion-in-Context Learning, a novel approach that enriches ICL by incorporating a more diverse set of in-context examples, leading to better performance than standard ICL.
4.4.2 Reasoning
Structured reasoning techniques, such as chain-of-thought (CoT), decompose complex reasoning into smaller sequential steps, enhancing coherence and robustness in multimodal RAG systems. RAGAR Khaliq et al. (2024) introduces Chain of RAG and Tree of RAG to iteratively refine fact-checking queries and explore branching reasoning paths for more robust evidence generation.
VisDoM Suri et al. (2024) integrates CoT with evidence curation to ensure logical and contextual accuracy, while SAM-RAG Zhai (2024) employs reasoning chains alongside multi-stage answer verification to enhance the relevance, utility, and support of generated responses. Meanwhile, LDRE Yang et al. (2024) leverages LLMs for divergent compositional reasoning, generating refined captions by incorporating dense captions and modification text.
4.4.3 Instruction Tuning
@@
Several works have fine-tuned or instruct-tuned generation components for specific applications. RA-BLIP Ding et al. (2024b) leverages the Q-Former architecture from InstructBLIP Dai et al. (2023) to extract visual features based on question instructions, while RAGPT Lang et al. (2025) employs a context-aware prompter to generate dynamic prompts from relevant instances. mR2AG Zhang et al. (2024f) uses instruction tuning with the mR2AG-IT dataset to train MLLMs to adaptively invoke retrieval, identify relevant evidence, and generate accurate answers for knowledge-based VQA tasks.
RagVL Chen et al. (2024d) employs instruction tuning to enhance the ranking capability of MLLMs, serving them as a re-ranker for filtering the top-k retrieved images. Jang et al. (2024) focus on distinguishing image differences to generate descriptive textual responses. MMed-RAG Xia et al. (2024a) applies preference fine-tuning to help models balance retrieved knowledge with internal reasoning.
To improve generation quality, MegaPairs Zhou et al. (2024a) and Surf Sun et al. (2024a) construct multimodal instruction-tuning datasets from prior LLM errors, while Rule Xia et al. (2024b) refines Med-LVLM through direct preference optimization to mitigate overreliance on retrieved contexts.
4.4.4 Source Attribution and Evidence Transparency
Ensuring source attribution in multimodal RAG systems is a key focus of recent research. MuRAR Zhu et al. (2025) integrates multimodal data, fetched by a source-based retriever, to refine LLM’s initial response, ensuring informativeness. VISA Ma et al. (2024b) uses large vision-language models to generate answers with visual source attribution by identifying and highlighting supporting evidence in retrieved document screenshots. Similarly, OMG-QA Nan et al. (2024) ensures transparency by prompting the LLM to explicitly cite evidence in generated responses.
@@ 4.5 Training Strategies
Training multimodal RAG models involves a multi-stage process to effectively handle cross-modal interactions Chen et al. (2022a). Pretraining establishes the foundation using large paired datasets to learn cross-modal relationships while fine-tuning adapts models to specific tasks by leveraging cross-modal attention Ye et al. (2019). For instance, REVEAL Hu et al. (2023) integrates multiple training objectives. Its pretraining phase optimizes Prefix Language Modeling Loss ($L_{\text{PrefixLM}}$), where the model predicts text continuations from a prefix and an image. Supporting losses include Contrastive Loss ($L_{\text{contra}}$) for aligning queries with pseudo-ground-truth knowledge, Disentangled Regularization Loss ($L_{\text{decor}}$) to improve embedding expressiveness, and Alignment Regularization Loss ($L_{\text{align}}$) to align query and knowledge embeddings. During fine-tuning, a cross-entropy objective trains the model for tasks like VQA and image captioning. Details of formulas for widely used RAG loss functions can be found in Appendix (§D).
4.5.1 Alignment
@@
Contrastive learning improves representation quality by pulling positive pairs closer and pushing negative pairs apart in the embedding space. A common objective is the InfoNCE loss van den Oord et al. (2019), which maximizes the mutual information between positive pairs while minimizing similarity to negatives. Several multimodal RAG models, such as VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a) and SAM-RAG Zhai (2024) utilize InfoNCE loss to improve retrieval-augmented generation. Furthermore, EchoSight Yan and Xie (2024) enhances retrieval accuracy by selecting visually similar yet contextually distinct negatives, while HACL Jiang et al. (2024) improves generation by introducing hallucinative captions as distractors. Similarly, UniRaG Zhi Lim et al. (2024) strengthens retrieval by incorporating hard negative documents, helping the model distinguish relevant contexts from noise.
The eCLIP loss Kumar and Marttinen (2024) extends contrastive training by integrating expert-annotated data and an auxiliary Mean Squared Error (MSE) loss to refine embedding quality. Mixup strategies further augment training by generating synthetic positive pairs, improving generalization in contrastive learning Kumar and Marttinen (2024). Dense2Sparse Nguyen et al. (2024) incorporates two unidirectional losses: an image-to-caption loss $\ell(I\to C)$ and a caption-to-image loss $\ell(C\to I)$. It enforces sparsity through L1 regularization, optimizing retrieval precision by balancing dense and sparse representations.
4.5.2 Generation
A key aspect of multimodal RAG is the generation ability. Autoregressive language models are typically trained using Cross-Entropy Loss Brown et al. (2020). For image generation, widely used approaches include Generative Adversarial Networks (GANs) Goodfellow et al. (2014) and Diffusion Models Ho et al. (2020). GANs employ various loss functions, such as Binary Cross-Entropy Loss, Minimax Loss, Wasserstein Loss (WGAN), and Hinge Loss. Diffusion Models utilize Mean Squared Error (MSE) Loss for noise prediction, a common approach in Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. (2020).
4.5.3 Robustness and Noise Management
Multimodal training faces challenges such as noise and modality-specific biases Buettner and Kovashka (2024). Managing noisy retrieval inputs is critical for maintaining model performance. MORE Cui et al. (2024) injects irrelevant results during training to enhance focus on relevant inputs. AlzheimerRAG Lahiri and Hu (2024) uses progressive knowledge distillation to reduce noise while maintaining multimodal alignment. RAGTrans Cheng et al. (2024) leverages hypergraph-based knowledge aggregation to refine multimodal representations, ensuring more effective propagation of relevant information. RA-BLIP Ding et al. (2024b) introduces the Adaptive Selection Knowledge Generation (ASKG) strategy, which leverages the implicit capabilities of LLMs to filter relevant knowledge for generation through a denoising-enhanced loss term, eliminating the need for fine-tuning. This approach achieves strong performance compared to baselines while significantly reducing computational overhead by minimizing trainable parameters.
RagVL Chen et al. (2024d) improves robustness through noise-injected training by adding hard negative samples at the data level and applying Gaussian noise with loss reweighting at the token level, enhancing the model’s resilience to multimodal noise. Finally, RA-CM3 Yasunaga et al. (2023) enhances generalization using Query Dropout, which randomly removes query tokens during retrieval, serving as a regularization method that improves generator performance.
5 Tasks Addressed by Multimodal RAGs
Multimodal RAG systems extend RAG beyond unimodal settings to tasks requiring cross-modal integration, enhancing performance across modalities such as text, images, and audio. In content generation, these models enhance image captioning Zhi Lim et al. (2024); Hu et al. (2023); Rao et al. (2024) and text-to-image synthesis Yasunaga et al. (2023); Chen et al. (2022b) by retrieving relevant contextual information. They also improve coherence in visual storytelling and ensure factual alignment in multimodal summarization Tonmoy et al. (2024). In knowledge-intensive applications, multimodal RAG supports open-domain and knowledge-seeking question answering Chen et al. (2024d); Ding et al. (2024b); Yuan et al. (2023), video-based QA Luo et al. (2024b), and fact verification Khaliq et al. (2024), grounding responses in retrieved knowledge and thereby mitigating hallucinations.
Cross-modal retrieval further advances zero-shot image–text retrieval Yang et al. (2024); Dong et al. (2024b). Additionally, the recent incorporation of chain-of-thought reasoning Zhai (2024); Khaliq et al. (2024) has enhanced its ability to support complex problem solving and logical inference. Finally, integrating multimodal RAG into interactive agents and AI assistants such as Gemini Team et al. (2024) enables natural language-driven visual search, document understanding, and multimodal reasoning. The taxonomy of application domains can be seen in Figure 3. The following sections explore domain-specific adaptations of these techniques in greater depth.
Healthcare and Medicine
Multimodal RAG enhances clinical decision-making through integrated analysis of medical imaging, electronic health records, and biomedical literature. Systems like MMED-RAG Xia et al. (2024a) address diagnostic uncertainty in medical visual question answering by aligning radiology images with contextual patient data. RULE Xia et al. (2024b) mitigates hallucinations in automated report generation through dynamic retrieval of clinically similar cases. AsthmaBot Bahaj and Ghogho (2024) introduces a multimodal RAG-based approach for supporting asthma patients across multiple languages, enabling structured, language-specific semantic searches. Predictive frameworks such as Realm Zhu et al. (2024c) demonstrate robust risk assessment by fusing heterogeneous patient data streams, while Hybrid RAG Su et al. (2024a) advances privacy-preserving architectures for federated clinical data integration. FactMM-RAG Sun et al. (2024b) automates radiology report drafting by retrieving biomarker correlations from medical ontologies, exemplifying the paradigm’s capacity to operationalize expert knowledge at scale.
Software Engineering
Fashion and E-Commerce
Cross-modal alignment drives advancements in product discovery and design automation. UniFashion Zhao et al. (2024) enables style-aware retrieval by jointly embedding garment images and textual descriptors, while Dang (2024) reduces search friction through multimodal query expansion. LLM4DESIGN Chen et al. (2024c) demonstrates architectural design automation by retrieving compliance constraints and environmental impact assessments, underscoring RAG’s adaptability to creative domains.
Entertainment and Social Computing
Emerging Applications

### v3

@@ 3.2 Fusion Mechanisms
Score Fusion and Alignment
@@
Models in this category utilize distinct strategies to align multimodal representations. Zhi Lim et al. (2024) convert text, tables, and images into a single textual format using a cross-encoder trained for relevance scoring. Sharifymoghaddam et al. (2024) introduce interleaved image–text pairs that vertically merge multiple few-shot images (as in LLaVA Liu et al. (2023a)), while aligning modalities via CLIP score fusion Hessel et al. (2021) and BLIP feature fusion Li et al. (2022). Wiki-LLaVA Caffagni et al. (2024), C3Net Zhang et al. (2024c), Riedler and Langer (2024), and MegaPairs Zhou et al. (2024a) embed images and queries into a shared CLIP space. In particular, MegaPairs Zhou et al. (2024a) scales this approach by integrating both CLIP-based and MLLM-based retrieval, fusing their scores to leverage complementary strengths, but at the cost of increased inference complexity. VISA Ma et al. (2024b) employs the Document Screenshot Embedding (DSE) model to align textual queries with visual document representations by encoding both into a shared embedding space. REVEAL Hu et al. (2023) injects retrieval scores into attention layers to minimize L2-norm differences between query and knowledge embeddings, and MA-LMM He et al. (2024) aligns video-text embeddings via a BLIP-inspired Query Transformer Li et al. (2022). LLM-RA Jian et al. (2024) concatenates text and visual embeddings into joint queries to reduce retrieval noise, while RA-BLIP Ding et al. (2024b) employs a 3-layer BERT-based adaptive fusion module to unify visual–textual semantics. Xue et al. (2024b) use a prototype-based embedding network Zheng et al. (2023) to map object-predicate pairs into a shared semantic space, aligning visual features with textual prototypes. Re-IMAGEN Chen et al. (2022b) balances creativity and entity fidelity in text-to-image synthesis via interleaved classifier-free guidance during diffusion sampling. To improve multimodal alignment, VISRAG Yu et al. (2024) applies position-weighted mean pooling over VLM hidden states, giving higher relevance to later tokens. In contrast, RAG-Driver Yuan et al. (2024) aligns visual and language embeddings through visual instruction tuning and an MLP projector.
Attention-Based Mechanisms
@@
Attention-based methods dynamically modulate cross-modal interactions to enable fine-tuned reasoning across tasks, balancing specificity and interpretability. Cross-attention is frequently used to integrate heterogeneous modalities, as in EMERGE Zhu et al. (2024b), MORE Cui et al. (2024), and AlzheimerRAG Lahiri and Hu (2024), though often requiring task-specific attention heads. RAMM Yuan et al. (2023) employs a dual-stream co-attention transformer, combining self-attention and cross-attention to fuse retrieved biomedical images/texts with input data. RAGTrans Cheng et al. (2024) applies user-aware attention to social media features. MV-Adapter Jin et al. (2024) introduces Cross-Modality Tying to align video-text embeddings by sharing latent factors, improving robustness but reducing granularity of modality-specific features. M2-RAAP Dong et al. (2024b) enhances fusion through an auxiliary caption-guided strategy that re-weights frames and text captions based on intra-modal similarity, then uses a mutual-guided alignment head to filter misaligned features via dot-product similarity and frame-to-token attention; however, this method is computationally intensive. Xu et al. (2024a) condition text generation on visual features using gated cross-attention, optimizing controllability but requiring aligned supervision, and Mu-RAG Chen et al. (2022a) employs intermediate cross-attention for open-domain QA. Kim et al. (2024) leverage cross-modal memory retrieval with pre-trained CLIP ViT-L/14 to map video-text pairs into a shared space, enabling dense captioning through the attention-based fusion of retrieved memories.
Unified Frameworks and Projections
@@
Unified frameworks and projection methods consolidate multimodal inputs into coherent representations. Su et al. (2024a) employ hierarchical cross-chains and late fusion for healthcare data, while IRAMIG Liu et al. (2024b) iteratively integrates multimodal results into unified knowledge representations, enhancing consistency but requiring multiple reasoning passes. M3DocRAG Cho et al. (2024) flattens multi-page documents into a single embedding tensor, and PDF-MVQA Ding et al. (2024d) proposes a joint-grained retriever that fuses coarse-grained semantic entity representations with their fine-grained token-level textual content, creating a richer, unified representation. DQU-CIR Wen et al. (2024) unifies raw data by converting images into text captions for complex queries and overlaying text onto images for simple ones, then fusing embeddings via MLP-learned weights. SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input to unimodal text for subsequent processing. UFineBench Zuo et al. (2024) uses a shared granularity decoder for ultra-fine text–person retrieval. Nguyen et al. (2024) introduce Dense2Sparse projection, converting dense embeddings from models like BLIP/ALBEF Li et al. (2022) into sparse lexical vectors using layer normalization and probabilistic expansion control to optimize storage and interpretability.
@@ 3.3 Augmentation Techniques
Basic RAG systems typically retrieve content in a single step, directly passing it to generation, often leading to inefficiencies and suboptimal outputs. Augmentation techniques refine retrieved data beforehand, improving multimodal interpretation, structuring, and integration Gao et al. (2023).
Context Enrichment
@@
This focuses on enhancing the relevance of retrieved knowledge by refining or expanding retrieved data. General approaches incorporate additional contextual elements (e.g., text chunks, image tokens, structured data) to provide a richer grounding for generation Caffagni et al. (2024); Xue et al. (2024b). EMERGE Zhu et al. (2024b) enriches context by integrating entity relationships and semantic descriptions. MiRAG Adjali et al. (2024) expands initial queries through entity retrieval and reformulation, enhancing subsequent stages for the visual question-answering. Video-RAG Luo et al. (2024b) enhances long-video understanding through Query Decoupling, which reformulates user queries into structured retrieval requests to extract auxiliary multimodal context. Img2Loc Zhou et al. (2024e) boosts accuracy by including both similar and dissimilar points in prompts, helping rule out implausible locations.
Adaptive and Iterative Retrieval
@@
For more complex queries, dynamic retrieval mechanisms have proven effective. Adaptive retrieval approaches optimize relevance by adjusting retrieval dynamically. For instance, UniversalRAG Yeo et al. (2025) introduces a framework that adapts retrieval by dynamically routing queries to the most suitable corpus based on both the required modality and granularity (e.g., paragraph vs. document, clip vs. full video), thereby addressing the specific knowledge type and scope demanded by the query. SKURG Yang et al. (2023) determines the number of retrieval hops based on query complexity. SAM-RAG Zhai (2024) and mR2AG Zhang et al. (2024f) dynamically assess the need for external knowledge and filter irrelevant content using MLLMs to retain only task-critical information. MMed-RAG Xia et al. (2024a) further improves retrieval precision by discarding low-relevance results, while OmniSearch Li et al. (2024d) decomposes multimodal queries into structured sub-questions, planning retrieval actions in real time. Iterative approaches refine results over multiple steps by incorporating feedback from prior iterations. For example, OMGM Yang et al. (2025) orchestrates a multi-step, coarse-to-fine retrieval process for knowledge-based visual question answering, starting with a broad entity search and progressively refining the selection through multimodal reranking and fine-grained textual filtering to pinpoint the most relevant knowledge, achieving superior retrieval performance in comparison to prior methods. IRAMIG Liu et al. (2024b) improves multimodal retrieval by dynamically updating queries based on retrieved content. OMG-QA Nan et al. (2024b) integrates episodic memory to refine retrieval across multiple rounds, ensuring continuity in reasoning. RAGAR Khaliq et al. (2024) further enhances contextual consistency by iteratively adjusting retrieval based on prior responses and multimodal analysis.
@@ 3.4 Generation Techniques
In-Context Learning (ICL)
@@
ICL with retrieval augmentation enhances reasoning in multimodal RAGs by leveraging retrieved content as few-shot examples without requiring retraining. Models such as RMR Tan et al. (2024), Sharifymoghaddam et al. (2024), and RA-CM3 Yasunaga et al. (2023), extend this paradigm to multimodal RAG settings. RAG-Driver Yuan et al. (2024) refines ICL by retrieving relevant driving experiences from a memory database. MSIER Luo et al. (2024a) improves example selection with a multimodal supervised in-context examples retrieval framework, using an MLLM scorer to assess textual and visual relevance. Raven Rao et al. (2024) introduces Fusion-in-Context Learning, integrating diverse in-context examples for superior performance over standard ICL.
Reasoning
Reasoning methods, like chain of thought (CoT), decompose complex reasoning into sequential steps, improving coherence and robustness in multimodal RAG. RAGAR Khaliq et al. (2024) refines fact-checking queries and explores branching reasoning paths by introducing Chain of RAG and Tree of RAG, while VisDoM Suri et al. (2024) and SAM-RAG Zhai (2024) integrate CoT with evidence curation and multi-stage verification to enhance accuracy and support. Notably, VisDoM performs well in scenarios where key information is distributed across modalities. LDRE Yang et al. (2024b) applies LLMs for divergent compositional reasoning by refining captions using dense descriptions and textual modifications, achieving superior zero-shot composed image retrieval results.
Instruction Tuning
@@
Several works have fine-tuned or instruct-tuned generation components for specific applications. RA-BLIP Ding et al. (2024b) leverages the Q-Former architecture from InstructBLIP Dai et al. (2023) to extract visual features based on question instructions, while RAGPT Lang et al. (2025) employs a context-aware prompter to generate dynamic prompts from relevant instances. MR2AG Zhang et al. (2024f) and RagVL Chen et al. (2024e) train MLLMs to invoke retrieval adaptively, identify relevant evidence, and enhance ranking capabilities for improved response accuracy. Jang et al. (2024) focus on distinguishing image differences to generate descriptive textual responses. MMed-RAG Xia et al. (2024a) applies preference fine-tuning to help models balance retrieved knowledge with internal reasoning. To improve generation quality, MegaPairs Zhou et al. (2024a) and Surf Sun et al. (2024a) construct multimodal instruction-tuning datasets from prior LLM errors, while Rule Xia et al. (2024b) refines a medical large vision language model through direct preference optimization to mitigate overreliance on retrieved contexts.
Source Attribution and Evidence Transparency
Ensuring source attribution in multimodal RAG systems is a significant research focus. OMG-QA Nan et al. (2024b) prompts LLMs for explicit evidence citation in generated responses. MuRAR Zhu et al. (2025) refines an LLM’s initial response by integrating multimodal information from a source-based retriever to improve informativeness. However, its recall is constrained, as the retriever may miss evidence spanning different sections or web documents. Similarly, VISA Ma et al. (2024b) employs vision-language models to generate answers with visual source attribution by highlighting evidence in retrieved screenshots. Nevertheless, its attribution accuracy degrades when evidence spans multiple sections or requires cross-modal integration.
Agentic Generation and Interaction
Agent-driven multimodal RAG uses versatile autonomous/semi-autonomous systems across diverse interaction paradigms and specialized domains, often generating complex outputs. For user interaction, AppAgent v2 Li et al. (2024c) enables mobile GUI navigation while USER-LLM R1 Rahimi et al. (2025) creates personalized conversational agents via dynamic profiling, particularly for elderly users. In specialized applications, MMAD Jiang et al. (2025) addresses industrial anomaly detection with training-free enhancement strategies, Yi et al. (2025) improve clinical report generation while reducing hallucination, and CollEX Schneider et al. (2025) facilitates scientific collection exploration for researchers and learners. For complex reasoning, HM-RAG Liu et al. (2025a) coordinates hierarchical multi-agent collaboration across multimodal data streams, while CogPlanner Yu et al. (2025) introduces a cognitively inspired planning framework that iteratively refines queries and selects retrieval strategies adaptively.
@@ 3.5 Training Strategies
Training multimodal RAG models follows a multi-stage process to effectively capture cross-modal interactions Chen et al. (2022a). Pretraining on large paired datasets establishes cross-modal relationships, while fine-tuning adapts models to task-specific objectives by aligning outputs with task requirements Ye et al. (2019). For example, REVEAL Hu et al. (2023) integrates multiple training objectives. Its pretraining phase optimizes Prefix Language Modeling Loss ($L_{\text{PrefixLM}}$), where text is predicted from a given prefix and an associated image. Supporting losses include Contrastive Loss ($L_{\text{contra}}$) which aligns queries with pseudo-ground-truth knowledge, Disentangled Regularization Loss ($L_{\text{decor}}$) to enhance embedding expressiveness, and Alignment Regularization Loss ($L_{\text{align}}$) to refine query-knowledge alignment. Fine-tuning employs a cross-entropy objective for downstream tasks like visual question answering or image captioning. Details on robustness advancements and loss formulations are in Appendix (§D).
Alignment
@@
Contrastive learning improves representation quality by pulling positive pairs closer and pushing negative pairs apart in the embedding space. The InfoNCE loss van den Oord et al. (2019) is widely employed in multimodal RAG models, including VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a), and SAM-RAG Zhai (2024), to improve retrieval-augmented generation. Several models introduce refinements to contrastive training. EchoSight Yan and Xie (2024) enhances retrieval accuracy by selecting visually similar yet semantically distinct negatives, while HACL Jiang et al. (2024) mitigates hallucinations by incorporating adversarial captions as distractors. Similarly, UniRaG Zhi Lim et al. (2024) improves retrieval robustness by leveraging hard negative documents to help the model discriminate between relevant and irrelevant contexts. The eCLIP loss Kumar and Marttinen (2024) extends contrastive learning by integrating expert-annotated data and an auxiliary MSE loss to refine embedding quality. Mixup strategies further improve generalization by generating synthetic positive pairs Kumar and Marttinen (2024). Dense2Sparse Nguyen et al. (2024) employs image-to-caption $\ell(I\to C)$ and caption-to-image $\ell(C\to I)$ losses, while enforcing sparsity through $\ell 1$ regularization, optimizing retrieval precision by balancing dense and sparse representations.